OpenStudy (anonymous):

Solve 8.1x – 3.2 = 7.6x – 2.4. Answer x = 0.4 x = 1.6 This is an identity and is true for all real numbers. There is no solution.

OpenStudy (anonymous):

x=1.6

OpenStudy (anonymous):

x=16

OpenStudy (anonymous):

8.1x-7.6x=-2.4+3.2 0.5x=0.8 x=1.6

OpenStudy (waleed_imtiaz):

surely... x=1.6...... simplification